Making monthly payments, What will you owe at the end of 10 years if you took a loan of $125,000 from a bank for 12 years at an APR of 4.5% per year compounded daily if you paid the bank your entire bonus of $12,000 at the end of 5 years?
Answer #1

rate compounded monthly=((1+4.5%/365)^(365/12)-1)*12=4.5081696%

Monthly payment=125000*(4.5081696%/12)/(1-1/(1+4.5081696%/12)^(12*12))=1125.51436273

Loan outstanding after 5 years after paying bonus=125000*(1+4.5081696%/12)^(12*5)-1125.51436273/(4.5081696%/12)*((1+4.5081696%/12)^(12*5)-1)-12000=68949.18672358

Loan owed after 10 years=68949.18672358*(1+4.5081696%/12)^(12*5)-1125.51436273/(4.5081696%/12)*((1+4.5081696%/12)^(12*5)-1)=10756.45362750